Problem 12-23 Make or Buy Decision [LO12-3] Silven Industries, which manufactures and sells a highly successful line of summe5. What is the maximum price that Silven should be willing to pay the outside supplier for a box of 24 tubes? 6. Instead of sReq 7 Req 5 Req 6 Req 4 Req 3 Req 2 Req 1 What is the financial advantage (disadvantage) in total (not per box) if Silven buyReq 6 Req 7 Req 4 Req 5 Req 3 Req 2 Req 1 Should Silven Industries make or buy the tubes? OMake OBuyReq 2 Req 3 Req 5 Req 6 Req 7 Req 1 Req 4 Instead of sales of 155,000 boxes of tubes, revised estimates show a sales volume o
Answer #1

Solution 1:

Total manufacturing overhead per unit at 155000 boxes = $1.90 per unit

Fixed manufacturing overhead per unit = $139,500 / 155000 = $0.90 per unit

Variable manufacturing overhead per unit = $1.90 - $0.90 = $1 per unit

Reduction in Chap-Off manufacturing costs per box, if tubes purchased from outside vendor = (Direct labor cost per unit + Variable manufacturing overhead per unit) * 10% + Direct material cost per unit * 25%

= ($2.60 + $1)*10% + ($4.30*20%) = $1.22 per unit

Solution 2:

financial advantage (disadvantage) per box of Chap-Off if Silven buys its tubes from the outside supplier = $1.22 - $1.35 = ($0.13)

Solution 3:

Financial advantage (disadvantage) in total (not per box) if Silven buys 155,000 boxes of tubes from the outside supplier = -$0.13 * 155000 = ($20,150)

Solution 4:

As there is net financial disadvantage of $20,150, therefore silven should make the tubes.

Solution 5:

Maximum price that silven willing to pay outside supplier = Variable cost to make = $1.22 per box

Solution 6:

Cost to make 191000 boxes = (191000*$1.22) + $56,000 = $289,020

Cost to buy 191000 boxes = 191000*$1.35 = $257,850

Financial advantage (Disadvantage) of buying = $289,020 - $257,850 = $31,170

As there is financial advantage, therefore silven should buy the boxes.

Solution 7:

If outside supplier can accept any quantity of order, then silven should make 155000 boxes and buy 36000 boxes from outside supplier.